can someone help with solving this problem?
cubed root(x^2) = 4

Answer by mukhopadhyay(490)

cubed root(x^2) = 4
=> (x^2)^1/3 = 4
=> [(x^2)^1/3]^3 = 4^3
=> x^2 = 64
=> x = 8 (x=-8 ruled out because cubed root(x^2) has been specified as +ve)